# Junhui Hou

### v2

@@ Abstract
Cross-modal contrastive distillation has recently been explored for learning effective 3D representations. However, existing methods focus primarily on modality-shared features, neglecting the modality-specific features during the pre-training process, which leads to suboptimal representations. In this paper, we theoretically analyze the limitations of current contrastive methods for 3D representation learning and propose a new framework, namely CMCR, to address these shortcomings. Our approach improves upon traditional methods by better integrating both modality-shared and modality-specific features. Specifically, we introduce masked image modeling and occupancy estimation tasks to guide the network in learning more comprehensive modality-specific features. Furthermore, we propose a novel multi-modal unified codebook that learns an embedding space shared across different modalities. Besides, we introduce geometry-enhanced masked image modeling to further boost 3D representation learning. Extensive experiments demonstrate that our method mitigates the challenges faced by traditional approaches and consistently outperforms existing image-to-LiDAR contrastive distillation methods in downstream tasks. Code will be available at https://github.com/Eaphan/CMCR.
Keywords:
3D self-supervised learning Contrastive learning Vector-quantization 3D scene understanding
@@ 1 Introduction
A widely used approach for learning 3D representations is contrastive pixel-to-point knowledge transfer, which leverages synchronized and calibrated images and point clouds (Sautier et al., 2022; Mahmoud et al., 2023; Liu et al., 2024; Chen et al., 2024; Liao et al., 2024; Puy et al., 2024; Xu et al., 2025). The PPKT method (Liu et al., 2021) allows a 3D model to benefit from the extensive knowledge encoded in a pre-trained 2D image backbone by using a pixel-to-point contrastive loss, with no need for labeled data for either the images or point clouds. Following this, SLidR (Sautier et al., 2022) introduces superpixels to group pixels and points that come from visually coherent regions, resulting in a more structured contrastive task. Building on these ideas, Seal (Liu et al., 2024) leverages semantically rich superpixels generated by visual foundation models, incorporating temporal consistency regularization to enforce stability across point segments over time. Furthermore, CSC (Chen et al., 2024) investigates cross-scene semantic consistency for multi-modal 3D pre-training, aiming to ensure semantic coherence across all frames and scenes.
While contrastive learning methods have shown success in transferring knowledge between 2D and 3D modalities, they primarily focus on modality-shared information, which can limit their ability to fully capture the unique characteristics of each modality. These approaches emphasize aligning shared features across different modalities, but they often overlook the modality-specific details that could provide complementary insights. For instance, 3D point clouds contain rich spatial and geometric information, while 2D images encode fine-grained visual textures and color details. By concentrating on shared representations, these methods may miss out on leveraging the full potential of modality-specific features during pre-training, leading to suboptimal performance in downstream tasks that require a deeper understanding of each modality’s unique contributions (Xu et al., 2013; Liang et al., 2024).
@@ 2 Related Work
3D Scene Understanding. Traditional methods for 3D scene understanding often rely on representations such as raw points (Choe et al., 2022; Chen et al., 2022a), voxels (Puy et al., 2023), range views (Kong et al., 2023a; Tian et al., 2022), and multi-view fusion (Fadadu et al., 2022; Xu et al., 2021) to capture environmental features. While effective, these approaches typically depend on large volumes of labeled data, which is both time-consuming and expensive to obtain, thereby hindering the scalability of 3D perception models (Liu et al., 2022b). To mitigate this issue, recent research has explored alternatives to reduce the reliance on fully annotated datasets. These include semi-supervised (Kong et al., 2023b; Ho et al., 2024), weakly-supervised (Liu et al., 2023; Chibane et al., 2022), self-supervised, and active learning (Luo et al., 2023; Xie et al., 2023) strategies, all of which aim to leverage minimal labeled data or automatically generate useful annotations to enhance model performance and scalability.
@@
3D Representation Learning. Recent advancements in self-supervised learning for 3D point clouds have evolved alongside improvements in image-based methods (Zhang et al., 2024). These include pretext tasks like predicting transformations or reconstructing point cloud parts (Poursaeed et al., 2020; Sauder and Sievers, 2019). Discriminative methods focus on contrastive learning across different levels of representation (point, segment, region) to capture geometric and structural information (Nunes et al., 2022; Yin et al., 2022; Chen et al., 2022b). Temporal-consistency methods leverage spatiotemporal correlations, aligning objects over time for robust representations (Nunes et al., 2023; Huang et al., 2021). Reconstruction-based methods, such as those using Chamfer distance or surface reconstruction, aim to recover point cloud details from masked data. Cross-modal distillation approaches utilize synchronized camera-LiDAR data for contrastive learning, transferring knowledge from 2D to 3D networks (Sautier et al., 2022; Mahmoud et al., 2023; Liu et al., 2024; Chen et al., 2024; Liao et al., 2024; Puy et al., 2024; Xu et al., 2025). In this paper, we identify the limitations of these cross-distillation methods theoretically and propose a novel framework for more comprehensive 3D representation learning.
@@
Vector-Quantization and Codebook. Vector quantization (VQ) is a technique originally introduced for image generation, where a large set of vectors is partitioned into clusters, each represented by a code vector from a codebook (Van Den Oord et al., 2017; Peng et al., 2022). VQ was integrated into the autoencoder framework as VQ-VAE (Van Den Oord et al., 2017), where it enables discrete latent representations by converting an image into a sequence of discrete codes and reconstructing it from these codes. This approach not only facilitates more compact and stable latent representations but also addresses issues like posterior collapse and variance instability that often affect traditional VAEs. This technique has since been widely applied in various domains, including multimodal learning. Recent works have explored how to achieve a unified representation across multiple modalities by using a shared codebook (Liu et al., 2022a; Xia et al., 2024). For example, Liu et al. (Liu et al., 2022a) proposed using a unified discrete space for aligning short videos and speech/text, addressing the codebook cold-start problem with a code warm-up strategy. Chen et al. (2023) introduced FDT, which uses differentiable operations and can be trained end-to-end. In this work, we design a new multi-modal unified codebook to prevent the model from partitioning the codebook into modality-specific subspaces. Further technical details of our approach will be discussed in Section 4.2.
@@ 4.1 Overview
As depicted in Fig. 3, we propose a new 3D self-supervised learning method, namely CMCR (Cross-Modal Comprehensive Representation Learning), based on our deduction in Sec. 3. For decoupling two types of features, we develop heads $\{h_{\mathrm{2D}}^{\mathrm{sh}},h_{\mathrm{3D}}^{\mathrm{sh}}\}$ and $\{h_{\mathrm{2D}}^{\mathrm{sp}},h_{\mathrm{3D}}^{\mathrm{sp}}\}$ to extract the modality-shared features $\{F^{\mathrm{3D}},F^{\mathrm{2D}}\}$ and modality-specific features $\{G^{\mathrm{3D}},G^{\mathrm{2D}}\}$, respectively. We perform contrastive learning based on the modality-shared features of point-pixel pairs. Then we perform vector quantization on these modality-shared features based on a multi-modal unified codebook module. Next, those embedding vectors and the modality-specific features are added together for masked image restoration and occupancy estimation. These reconstruction tasks encourage the network to learn not only modality-shared features but also modality-specific features.
@@ 5.2 Transfer on 3D Semantic Segmentation
Quantitative Results. Under the linear probing setting, our method achieves the highest mIoU of 51.23% on nuScenes, significantly outperforming the previous state-of-the-art method, CSC (Chen et al., 2024), which records a mIoU of 46.00% (see Table 2). This highlights the superior representation quality of our pre-trained features without task-specific fine-tuning. For fine-tuning on nuScenes, our method consistently excels, particularly in low-data regimes, achieving a mIoU of 51.76% with just 1% of annotations, significantly higher than CSC (47.00%) and Seal (Liu et al., 2024) (45.84%). This trend persists across other proportions, with our method attaining 61.08% mIoU at 5% data and 65.69% at 10%, consistently outperforming all baselines. At 100% annotation, our method achieves a mIoU of 76.34%, surpassing CSC (75.70%) and Seal (75.60%). On SemanticKITTI, with only 1% annotation, our method achieves 49.86% mIoU, outperforming CSC (47.20%) and Seal (46.63%). These results demonstrate the robustness and versatility of our approach across diverse datasets and annotation settings.
@@ 5.4 Transfer on 3D Panoptic Segmentation
As shown in Table 6, our method consistently outperforms the others across all evaluation metrics: Panoptic Quality (PQ), Segmentation Quality (SQ), and Recognition Quality (RQ). At the 1% label setting, our approach achieves the highest scores in all three metrics, with a PQ of 20.7, SQ of 80.5, and RQ of 26.1, surpassing the second-best model (CSC) by 1.4% PQ, 6.0% SQ, and 1.5% RQ. This performance improvement remains consistent with 5% labeled data, where our method achieves a PQ of 24.0%, SQ of 78.5%, and RQ of 29.3%, outpacing CSC again by 0.9% PQ, 1.6% SQ, and 0.8% RQ.
@@ 6 Conclusion
In this paper, we have presented a novel framework for learning more comprehensive 3D representations. By addressing the limitations of existing methods that focus primarily on modality-shared features, we introduced a new approach that also captures modality-specific features through masked image modeling and occupancy estimation tasks. Our key contribution, the multi-modal unified codebook, enables the learning of shared embedding space across different modalities, facilitating better cross-modal alignment and representation learning. Moreover, the geometry-enhanced masked image modeling further enhances 3D representation learning by incorporating spatial structure information. Through extensive experiments, we demonstrated that our approach significantly improves upon traditional methods and outperforms existing image-to-LiDAR contrastive distillation methods in downstream tasks. Future work could explore additional task-specific adaptations and further optimizations to improve the performance of multi-modal 3D learning.
@@ Acknowledgements
This work was supported in part by the NSFC Excellent Young Scientists Fund 62422118, and in part by the Hong Kong Research Grants Council under Grants 11219324 and 11219422.

### v3

@@ Abstract
Cross-modal contrastive distillation has recently been explored for learning effective 3D representations. However, existing methods focus primarily on modality-shared features, neglecting the modality-specific features during the pre-training process, which leads to suboptimal representations. In this paper, we theoretically analyze the limitations of current contrastive methods for 3D representation learning and propose a new framework, namely CMCR (Cross-Modal Comprehensive Representation Learning), to address these shortcomings. Our approach improves upon traditional methods by better integrating both modality-shared and modality-specific features. Specifically, we introduce masked image modeling and occupancy estimation tasks to guide the network in learning more comprehensive modality-specific features. Furthermore, we propose a novel multi-modal unified codebook that learns an embedding space shared across different modalities. Besides, we introduce geometry-enhanced masked image modeling to further boost 3D representation learning. Extensive experiments demonstrate that our method mitigates the challenges faced by traditional approaches and consistently outperforms existing image-to-LiDAR contrastive distillation methods in downstream tasks. Code will be available at github.com/Eaphan/CMCR.
@@ 1 Introduction
A widely used approach for learning 3D representations is contrastive pixel-to-point knowledge transfer, which leverages synchronized and calibrated images and point clouds (Sautier et al., 2022; Mahmoud et al., 2023; Liu et al., 2023b; Chen et al., 2024; Liao et al., 2024; Puy et al., 2024; Xu et al., 2025). The PPKT method (Liu et al., 2021) allows a 3D model to benefit from the extensive knowledge encoded in a pre-trained 2D image backbone by using a pixel-to-point contrastive loss, with no need for labeled data for either the images or point clouds. Following this, SLidR (Sautier et al., 2022) introduces superpixels to group pixels and points that come from visually coherent regions, resulting in a more structured contrastive task. Building on these ideas, Seal (Liu et al., 2023b) leverages semantically rich superpixels generated by visual foundation models, incorporating temporal consistency regularization to enforce stability across point segments over time. Furthermore, CSC (Chen et al., 2024) investigates cross-scene semantic consistency for multi-modal 3D pre-training, aiming to ensure semantic coherence across all frames and scenes.
While contrastive learning methods have shown success in transferring knowledge between 2D and 3D modalities, they primarily focus on modality-shared information, which can limit their ability to fully capture the unique characteristics of each modality. These approaches emphasize aligning shared features across different modalities, but they often overlook the modality-specific details that could provide complementary insights. For instance, 3D point clouds contain rich spatial and geometric information, while 2D images encode fine-grained visual textures and color details. By concentrating on shared representations, these methods may miss out on leveraging the full potential of modality-specific features during pre-training, leading to suboptimal performance in downstream tasks that require a deeper understanding of each modality’s unique contributions (Xu et al., 2013; Liang et al., 2023).
@@ 2 Related Work
3D Scene Understanding. Traditional methods for 3D scene understanding often rely on representations such as raw points (Choe et al., 2022; Chen et al., 2022a), voxels (Puy et al., 2023), range views (Kong et al., 2023a; Tian et al., 2022), and multi-view fusion (Fadadu et al., 2022; Xu et al., 2021) to capture environmental features. While effective, these approaches typically depend on large volumes of labeled data, which is both time-consuming and expensive to obtain, thereby hindering the scalability of 3D perception models (Liu et al., 2022b). To mitigate this issue, recent research has explored alternatives to reduce the reliance on fully annotated datasets. These include semi-supervised (Kong et al., 2023b; Ho et al., 2023), weakly-supervised (Liu et al., 2023a; Chibane et al., 2022), self-supervised, and active learning (Luo et al., 2023; Xie et al., 2023) strategies, all of which aim to leverage minimal labeled data or automatically generate useful annotations to enhance model performance and scalability.
@@
3D Representation Learning. Recent advancements in self-supervised learning for 3D point clouds have evolved alongside improvements in image-based methods (Zhang et al., 2024). These include pretext tasks like predicting transformations or reconstructing point cloud parts (Poursaeed et al., 2020; Sauder and Sievers, 2019). Discriminative methods focus on contrastive learning across different levels of representation (point, segment, region) to capture geometric and structural information (Nunes et al., 2022; Yin et al., 2022; Chen et al., 2022b). Temporal-consistency methods leverage spatiotemporal correlations, aligning objects over time for robust representations (Nunes et al., 2023; Huang et al., 2021). Reconstruction-based methods, such as those using Chamfer distance or surface reconstruction, aim to recover point cloud details from masked data. There is another type of method that conducts pre-training through cross-modal knowledge distillation, as detailed below.
Image-to-LiDAR Contrastive Distillation. Cross-modal distillation approaches utilize synchronized camera-LiDAR data for contrastive learning, transferring knowledge from 2D to 3D networks (Sautier et al., 2022; Mahmoud et al., 2023; Zhang and Hou, 2024; Chen et al., 2024; Liao et al., 2024; Puy et al., 2024; Xu et al., 2025).  Sautier et al. (2022) introduced a contrastive loss for self-supervised 2D-to-3D representation learning, using a superpixel-to-superpoint framework. This method was later expanded by Mahmoud et al. (2023), who integrated a semantically tolerant contrastive constraint along with a class-balancing loss to improve its robustness. Building on these advancements, Liu et al. (2023b) further refined the approach by incorporating spatial and temporal consistency regularization, which enhanced semantic feature learning. Additionally,  Zhang et al. (2024) extended cross-modality contrastive distillation beyond the image plane, exploring its application in bird-eye views to improve representation alignment across modalities. OLIVINE and CSC enhance 3D representation with VFM-guided semantic prototypes (Zhang and Hou, 2024; Chen et al., 2024). SuperFlow is a 4D spatio-temporal framework for LiDAR-camera contrastive pretraining that uses dense-to-sparse regularization and flow-based contrast to enhance 3D representations (Xu et al., 2025). In this paper, we identify the limitations of these cross-distillation methods theoretically and propose a novel framework for more comprehensive 3D representation learning.
@@
Vector-Quantization and Codebook. Vector quantization (VQ) is a technique originally introduced for image generation, where a large set of vectors is partitioned into clusters, each represented by a code vector from a codebook (Van Den Oord et al., 2017; Peng et al., 2022). VQ was integrated into the autoencoder framework as VQ-VAE (Van Den Oord et al., 2017), where it enables discrete latent representations by converting an image into a sequence of discrete codes and reconstructing it from these codes. This approach not only facilitates more compact and stable latent representations but also addresses issues like posterior collapse and variance instability that often affect traditional VAEs. This technique has since been widely applied in various domains, including multimodal learning. Recent works have explored how to achieve a unified representation across multiple modalities by using a shared codebook (Liu et al., 2022a; Xia et al., 2023). For example, Liu et al. (Liu et al., 2022a) proposed using a unified discrete space for aligning short videos and speech/text, addressing the codebook cold-start problem with a code warm-up strategy. Chen et al. (2023) introduced FDT, which uses differentiable operations and can be trained end-to-end. In this work, we design a new multi-modal unified codebook to prevent the model from partitioning the codebook into modality-specific subspaces. Further technical details of our approach will be discussed in Section 4.2.
@@ 4.1 Overview
Based on our deduction in Sec. 3, we propose a new 3D self-supervised learning method, namely CMCR (Cross-Modal Comprehensive Representation Learning), as depicted in Fig. 3. For decoupling two types of features, we develop heads $\{h_{\mathrm{2D}}^{\mathrm{sh}},h_{\mathrm{3D}}^{\mathrm{sh}}\}$ and $\{h_{\mathrm{2D}}^{\mathrm{sp}},h_{\mathrm{3D}}^{\mathrm{sp}}\}$ to extract the modality-shared features $\{F^{\mathrm{3D}},F^{\mathrm{2D}}\}$ and modality-specific features $\{G^{\mathrm{3D}},G^{\mathrm{2D}}\}$, respectively. We perform contrastive learning based on the modality-shared features of point-pixel pairs. Then we perform vector quantization on these modality-shared features based on a multi-modal unified codebook module. Next, those embedding vectors and the modality-specific features are added together for masked image restoration and occupancy estimation. These reconstruction tasks encourage the network to learn not only modality-shared features but also modality-specific features.
Choices of Modality-Specific Pretext Tasks. 1) For point clouds, we use occupancy estimation, which predicts if query points are inside or on the surface of an object. Unlike Masked Autoencoders (MAE), which face ambiguity in sparse point clouds, occupancy estimation provides a clear supervisory signal, enabling the model to learn 3D geometry and surface details. 2) MIM is selected for the image modality as it encourages the model to capture contextual information from visible regions, learning fine-grained texture, color, and spatial relationships unique to images. It also aligns 2D features with 3D representations for later fusion. Unlike Variational Autoencoders (VAEs), MIM does not require complex probabilistic modeling.
@@ 4.2 Multi-modal Unified Codebook
Remark. The multi-modal codebook provides structured semantic alignment between image and 3D point cloud features, enhancing cross-modal coherence. By mapping features to discrete codewords, the model captures both basic and complex shared patterns across modalities. Unlike InfoNCE-based methods, which only enforce feature similarity, the codebook offers a more sophisticated alignment, preserving semantic consistency and improving representation learning. This results in more robust and conceptually richer features beyond simple pairwise alignment.
@@ 5.2 Transfer on 3D Semantic Segmentation
Quantitative Results. Under the linear probing setting, our method achieves the highest mIoU of 51.23% on nuScenes, significantly outperforming the previous state-of-the-art method, CSC (Chen et al., 2024), which records a mIoU of 46.00% (see Table 2). This highlights the superior representation quality of our pre-trained features without task-specific fine-tuning. For fine-tuning on nuScenes, our method consistently excels, particularly in low-data regimes, achieving a mIoU of 51.76% with just 1% of annotations, significantly higher than CSC (47.00%) and Seal (Liu et al., 2023b) (45.84%). This trend persists across other proportions, with our method attaining 61.08% mIoU at 5% data and 65.69% at 10%, consistently outperforming all baselines. At 100% annotation, our method achieves a mIoU of 76.34%, surpassing CSC (75.70%) and Seal (75.60%). On SemanticKITTI, with only 1% annotation, our method achieves 49.86% mIoU, outperforming CSC (47.20%) and Seal (46.63%). These results demonstrate the robustness and versatility of our approach across diverse datasets and annotation settings.
@@ 5.4 Transfer on 3D Panoptic Segmentation
As shown in Table 6, our method consistently outperforms the others across all evaluation metrics: Panoptic Quality (PQ), Segmentation Quality (SQ), and Recognition Quality (RQ). At the 1% label setting, our approach achieves the highest scores in all three metrics, with a PQ of 20.7, SQ of 80.5, and RQ of 26.1, surpassing the second-best model (SuperFlow) by 0.9% PQ, 2.2% SQ, and 1.0% RQ. This performance improvement remains consistent with 5% labeled data, where our method achieves a PQ of 24.0%, SQ of 78.5%, and RQ of 29.3%, outpacing SuperFlow again by 0.6% PQ, 0.9% SQ, and 0.5% RQ.
@@ 6 Conclusion
In this paper, we have presented a novel framework for learning more comprehensive 3D representations. By addressing the limitations of existing methods that focus primarily on modality-shared features, we introduced a new approach that also captures modality-specific features through masked image modeling and occupancy estimation tasks. Our key contribution, the multi-modal unified codebook, enables the learning of a shared embedding space across different modalities, facilitating better cross-modal alignment and representation learning. Moreover, the geometry-enhanced masked image modeling further enhances 3D representation learning by incorporating spatial structure information. Through extensive experiments, we demonstrated that our approach significantly improves upon traditional methods and outperforms existing image-to-LiDAR contrastive distillation methods in downstream tasks. Future work could explore additional task-specific adaptations and further optimizations to improve the performance of multi-modal 3D learning.
@@ Acknowledgements
This work was supported in part by the National Natural Science Foundation of China Excellent Young Scientists Fund (Grant 62422118); in part by the Hong Kong Research Grants Council (Grants 11219324 and 11219422); and in part by the Shanghai Pujiang Program (Grant 25PJA042). It was also supported by Shanghai Technical Service Center of Science and Engineering Computing, Shanghai University.